# Post-processing improves accuracy of Artificial Intelligence weather forecasts

### v1

@@ 4 Results
As noted by Lang et al. (2024a) and Ben Bouallègue et al. (2024), ML forecasts tend to have a smoother appearance than physics-based NWP, because they are based on optimising a loss function, which rewards predicting the middle of the distribution, rather than simulating physical processes. Additionally, graph neural networks, including the graph transformer architecture used in AIFS, are susceptible to a problem where, as the number of layers increases, message-passing between nodes increases similarity between the features of different nodes (Wu et al., 2023), which would contribute to visual smoothness of the gridded output. Figure 3 compares gridded HRES and AIFS temperature forecasts before and after post-processing. The input AIFS is noticeably smoother and less detailed than HRES, although this is partly also due also to their different resolutions, 0.25°0.25°0.25  ° vs 0.1°0.1°0.1  °. [^2] For a comparison of AIFS vs the ECMWF integrated forecasting system (IFS) at the same resolution, see Figure 7 of Lang et al. (2024a). Similarly, Figure 3 in Ben Bouallègue et al. (2024) compares the IFS with Pangu-Weather at equal resolution. In both examples, the AI forecasts are smoother than their NWP counterparts. However, after post-processing, the outputs have a similar level of detail; in particular, it is worth noting that post-processing is able to add realistic topographic features.
@@
Figure 3: Raw forecasts (top row) and post-processed expected value outputs from IMPROVER (bottom row) for temperature at lead time 18H and valid time 2024-06-15 06:00 UTC (16:00 Australian Eastern Standard Time). The left column is HRES and the right AIFS. Units are degrees Celsius. Note that the raw forecasts are on a latitude/longitude grid, while the post-processed forecasts use the Albers equal area projection. The diagonal artefacts near the edges in the lower part of the plots on the right occur because calibration is done against the MSAS analysis, which has a more limited spatial domain.
@@
Figure 4 shows forecast mean squared error by lead day before and after IMPROVER post-processing. For the raw ENS forecast, the quantity verified is the ensemble mean. Lead day 0 corresponds to lead hours 0-24 inclusive, lead day 1 to hours 25-48, et cetera. Post-processing generally improves the raw forecasts, yielding gains of a few days of skill in some cases. For all variables, the three post-processed models achieve similar accuracy early in the forecast period. For temperature and dew point, after the first few lead days AIFS has a clear advantage over HRES, with the gap widening as lead time increases.
@@
It is well established that blending models generally improves the accuracy relative to the best individual forecast, and therefore it is a common approach in operational weather forecasts. Figure 5 compares the two different blends, along with the post-processed models. To measure the difference between the two blends, we used the Diebold-Mariano test to calculate a confidence interval for the difference in MSE for each lead day. This test requires that the time series of differences be weakly stationary. Therefore, in order to avoid diurnal and spatial biases in the error, for each lead time we aggregated the data by valid day and calculated the MSE (including all sites and valid times for that valid day). [^3] The aggregation is slightly different from the aggregation by lead day shown in Figure 4, which simply combines all data for each lead day, without the intermediate step of taking the mean by valid day. The test statistic was calculated on this daily series, using the implementation in the “scores” Python package (Leeuwenburg et al., 2024) based on the method of Hering and Genton (2011). Figure 6 shows the MSE difference with confidence interval. Overall, the blend including all three models is significantly better than the blend of only the two NWP models.
@@
Figure 4: Mean squared error by lead day for raw (dotted line) and post-processed (solid line) models ENS (blue), HRES (red) and AIFS (green), for temperature (left), dew point (middle) and wind speed (right). The calculation includes only lead times that are present in both the raw and post-processed forecasts (that is, those that are multiples of 6 hours).
@@
Figure 5: Mean squared error by lead day for post-processed models ENS (blue), HRES (red) and AIFS (green) and blends (NWP models ENS and HRES, yellow; and all models, brown). Note that the line for ENS is very close to that of the NWP-model blend, and partially hidden by it.
@@
Figure 7 shows the bias (defined as forecast minus observation) of the various models. As with MSE, for the raw ENS forecast, bias is calculated for the ensemble mean. Bias correction significantly mitigates the biases present in the raw forecasts. However some diurnal patterns and constant biases remain. This is to be expected because the IMPROVER bias correction is done relative to the MSAS analysis, which has its own biases relative to the observation data used for verification. We also observe trends in the bias over the forecast period. This is somewhat surprising, given that the bias correction and reliability calibration are calculated separately for each lead time. One possible explanation is that the calibration is based on a moving window of recent history, so at longer lead times there is a greater lag between the validity time of the forecast and the most recent ground truth used in the calibration. This perhaps reduces the ability of the calibration to correct for both seasonal model biases (that is, those varying systematically by time of year), and for idiosyncratic errors in the currently-occuring weather system.
@@
Figure 7: Bias of raw (dashed line) and post-processed (solid line) models ENS (blue), HRES (red) and AIFS (green) over the forecast period. Note that the raw models have 6-hour lead time frequency, while the post-processed outputs have 1-hour frequency.
@@
IMPROVER produces probabilistic outputs individually for both deterministic and ensemble input forecasts. For evaluating their accuracy, we use the continuous ranked probability score (CRPS), which measures the error between the predicted probability distribution and the true outcome. The CRPS is defined as the integral of the squared distance between the forecast distribution and the true outcome; formally, if $t:\to\mathrm{P}(X\leq t)$ is the predicted cumulative distribution function for some data point and $x$ the observed value, the CRPS for this observation is
[…]
@@
where $\mathbbm{1}_{(t\geq x)}$ is the indicator function which is equal to 1 if $t\geq x$ and 0 otherwise. Smaller CRPS values indicate a better forecast. Figure 8 shows the CRPS of the post-processed models and blends. While in general the blend of the two NWP models performs very similarly to the most accurate individual NWP model, the blend of all three models yields a larger improvment. We used the Diebold-Mariano test to compare the two blends, following the same process as described above for the MSE comparison. The results are shown in Figure 9; as is the case with MSE, we see that the all-model blend is significantly better than the NWP-model blend.
@@
Figure 8: Continuous rank probability score (CRPS) by lead day for post-processed models ENS (blue), HRES (red), and AIFS (green), and blends (NWP models ENS and HRES, yellow; and all models, brown) over the forecast period.
@@
Forecast reliability measuers how well-calibrated a forecast is. Specifically, for a given threshold $t$, a forecast has good reliability if the conditional probability $p_{o}$ that an observation exceeds $t$, given that the forecast probability is $p_{f}$, is approximately equal to $p_{f}$. Figure 10 shows the reliability at the 12-hour lead time (UTC 0) at selected thresholds, roughly corresponding to the 10th, 50th and 90th percentiles of the overall distribution of observations (including all times of day). There is no clear trend in these results: all models and blends have similar performance. The deterministic models arguably appear a little better than the ensemble, and the blended models do not seem to offer much advantage. It has been shown by Ranjan and Gneiting (2010) that blending probability forecasts does not preserve calibration; that is, even if each individual forecast is perfectly calibrated, in general the blend is not. Therefore, calibration could likely be further improved by recalibrating the blend. Functionality to implement the parametric recalibration proposed by Ranjan and Gneiting has recently been implemented in the IMPROVER software library, but is not currently part of the IMPROVER operational system used for this work. In Figure 11 we show the distributions of probabilistic forecasts at this lead time. The 3-model blend is in general less sharp than the 2-model blend, which is in turn less sharp than the sharpest NWP model. Again, this is expected for blended forecasts, and could be corrected with recalibration.
Figure 10: Reliability at the 12-hour lead time for post-processed models ENS (blue), HRES (red), and AIFS (green), and blends (NWP models ENS and HRES, yellow; and all models, brown). Forecast probabilities are binned into 10 equal-width bins, and, for each bin, the average forecast and observed probability of exceeding the threshold is calculated. These values are linearly interpolated to produce the reliability curve. Bins having fewer than 10 data points are not plotted.
Figure 11: Distribution of binned forecasts at the 12-hour lead time for post-processed models ENS (blue), HRES (red), and AIFS (green), and blends (NWP models ENS and HRES, yellow; and all models, brown). The $y$-axis uses the symmetric log scale, which is linear for values between 0 and 1, and logarithmic for larger values.
@@ 5 Discussion
This study demonstrates that established statistical post-processing methods, originally designed for traditional Numerical Weather Prediction (NWP) models, are also effective when applied to forecasts from the Artificial Intelligence Forecasting System (AIFS). Using the Australian Bureau of Meteorology’s operational IMPROVER system without modification, we observe comparable accuracy improvements in AIFS forecasts to those seen with NWP models. Notably, blending AIFS with traditional NWP models yields consistent gains in forecast skill for both expected value and probabilistic forecasts. These results indicate that AI-based forecasts can be integrated into operational systems using existing tools and workflows, thereby expanding the utility of these forecasts without the need for bespoke post-processing solutions. A primary finding of this work is that statistical calibration via IMPROVER substantially improves the accuracy of AIFS forecasts. The gains are comparable to those achieved when calibrating traditional models, and in some cases, calibration adds upwards of a day of forecast skill. This reinforces our view that AI models, like their physics-based counterparts, benefit from correction of systematic errors using historical forecast performance. Importantly, the post-processing workflow required no changes, suggesting a high degree of generality in these methods.
The calibrated AIFS output also produces high-quality probabilistic forecasts. For temperature and dew point in particular, the post-processed AIFS shows CRPS values similar to those of the ensemble system (ENS) during the early part of the forecast period. This is notable given that AIFS is a deterministic model, and the result highlights the value of fuzzy thresholding and reliability calibration in extracting probabilistic information from single-model inputs.
@@
Future work should consider extending these analyses to AIFS-CRPS (Lang et al., 2024b), an ensemble version of AIFS trained to minimise CRPS. This may further improve the skill of blended forecasts. Additionally, recalibrating the probabilistic blends – particularly in light of known issues with calibration preservation during blending – may yield further improvements.

### v2

@@ 4 Results
Our verification compares site-extracted forecast outputs with site observations, using the data described in Section 2. Note this is different to the ground truth used in the IMPROVER processing: since IMPROVER produces gridded forecasts, the bias correction and reliability calibration steps use the MSAS gridded analysis, as described in Section 33.1.
@@
As noted by Lang et al. (2024a) and Ben Bouallègue et al. (2024), ML forecasts tend to have a smoother appearance than physics-based NWP, because they are based on optimising a loss function, which rewards predicting the expected value of the distribution, rather than simulating physical processes. Additionally, graph neural networks, including the graph transformer architecture used in AIFS, are susceptible to a problem where, as the number of layers increases, message-passing between nodes increases similarity between the features of different nodes (Wu et al., 2023), which would contribute to visual smoothness of the gridded output. Figure 4 compares gridded HRES and AIFS temperature forecasts before and after post-processing. The input AIFS is noticeably smoother and less detailed than HRES, although this is partly also due also to their different resolutions, 0.25° vs 0.1°. [^7] For a comparison of AIFS vs the ECMWF integrated forecasting system (IFS) at the same resolution, see Figure 7 of Lang et al. (2024a). Similarly, Figure 3 in Ben Bouallègue et al. (2024) compares the IFS with Pangu-Weather at equal resolution. In both examples, the AI forecasts are smoother than their NWP counterparts. However, after post-processing, the AIFS forecast becomes more detailed and looks quite similar to HRES (although still a little smoother); in particular, post-processing is able to add realistic topographic features. Moreover, the fact that post-processed AIFS is slightly smoother than post-processed HRES does not seem to disadvantage it when assessing performance at sites: AIFS generally has better performance.
@@
Figure 4: Raw forecasts (top row) and post-processed expected value outputs from IMPROVER (bottom row) for temperature at lead time 18H and valid time 2024-06-15 06:00 UTC (16:00 Australian Eastern Standard Time). The left column is HRES and the right AIFS. Units are degrees Celsius. Note that the raw forecasts are on a latitude/longitude grid, while the post-processed forecasts use the Albers equal area projection. The diagonal artefacts near the edges in the post-processed forecasts occur because calibration is done against the MSAS analysis, which has a more limited spatial domain.
@@
Figure 5 shows forecast mean squared error by lead day before and after IMPROVER post-processing. For the raw ENS forecast, the quantity verified is the ensemble mean. Lead day 0 corresponds to lead hours 0-24 inclusive, lead day 1 to hours 25-48, et cetera. Post-processing generally improves the raw forecasts, yielding gains of a few days of skill in some cases. For all variables, the three post-processed models achieve similar accuracy early in the forecast period. For temperature and dew point, after the first few lead days AIFS has a clear advantage over HRES, with the gap widening as lead time increases; both are outperformed by ENS. The strong performance of AIFS is likely because it is trained specifically to minimise the MSE objective, rather than aiming to model the most likely forecast trajectory like HRES. However, as with NWP models, the autoregressive nature of AIFS means that small errors are amplified over the forecast period, which causes AIFS (and HRES) to fall behind ENS in accuracy as the lead time increases.
@@
It is well established that blending models generally improves the accuracy relative to the best individual forecast (see, for example, Vannitsem et al. (2021)), and therefore it is a common approach in operational weather forecasts. Figure 6 compares the two different blends, along with the post-processed models. To measure the difference between the two blends, we used the Diebold-Mariano test to calculate a confidence interval for the difference in MSE for each lead day. This test requires that the time series of differences be weakly stationary. Therefore, in order to avoid diurnal and spatial biases in the error, for each lead time we aggregated the data by valid day and calculated the MSE (including all sites and valid times for that valid day). [^8] The aggregation is slightly different from the aggregation by lead day shown in Figure 5, which simply combines all data for each lead day, without the intermediate step of taking the mean by valid day. The test statistic was calculated on this daily series, using the implementation in the “scores” Python package (Leeuwenburg et al., 2024) based on the method of Hering and Genton (2011). Figure 7 shows the MSE difference with confidence interval. Overall, the blend including all three models is significantly better than the blend of only the two NWP models.
To more closely examine the ways the different models contribute to the blend, we also considered two additional blends: a blend of HRES and AIFS, and a blend of ENS and AIFS. Figures 13 and 15 in the Appendix show the MSE and CRPS for the full array of blends (see also Tables 1, 2, 3 and Tables 4, 5, and 6 for the corresponding data). The blend of ENS and AIFS achieves very similar accuracy to the all-model blend. Interestingly, the blend of the only the two deterministic models, HRES and AIFS, also achieves similar accuracy to the all-model blend in the first few days of the forecast period. This offers the opportunity for operational centers to obtain high-quality short-term forecasts by using only deterministic forecasts, which are computionally much cheaper than ensembles. At later lead times the blends including ENS perform much more strongly, which is expected as the chaotic behaviour of the atmosphere makes deterministic forecasting difficult at these horizons.
@@
Figure 5: Mean squared error by lead day for raw (dashed line) and post-processed (solid line) models ENS (blue), HRES (red) and AIFS (green), for temperature (left), dew point (middle) and wind speed (right). The calculation includes only lead times that are present in both the raw and post-processed forecasts (that is, those that are multiples of 6 hours).
@@
Figure 6: Mean squared error by lead day for post-processed models ENS (blue dashed), HRES (red dashed) and AIFS (green dashed) and blends (NWP models ENS and HRES, yellow; and all models, pink). Note that the line for ENS is very close to that of the NWP-model blend, and partially hidden by it. The tabular data for these graphs can be found in Tables 1, 2 and 3 in the Appendix.
@@
Figure 8 shows the bias (defined as forecast minus observation) of the various models at the observation sites. As with MSE, for the raw ENS forecast, bias is calculated for the ensemble mean. Bias correction significantly mitigates the biases present in the raw forecasts. However some diurnal patterns and constant biases remain. This is to be expected because the IMPROVER bias correction is done relative to the MSAS analysis, which has its own biases relative to the site observation data used for verification (see Figure 14 in the Appendix). We also observe trends in the bias over the forecast period. This is somewhat surprising, given that the bias correction and reliability calibration are calculated separately for each lead time. One possible explanation is that the calibration is based on a moving window of the past 30 days’ history, so at longer lead times there is a greater lag between the validity time of the forecast and the most recent ground truth used in the calibration. This perhaps reduces the ability of the calibration to correct for both seasonal model biases (that is, those varying systematically by time of year), and for idiosyncratic errors in the currently-occuring weather system. An alternative approach could be to use a longer historical calibration dataset, containing a year or more of data, and calculate seasonally-varying biases. However, this requires a long history of model data, which is not available for new models like AIFS, and would also be more complex to implement operationally.
@@
Figure 8: Bias of raw (dotted line) and post-processed (solid line) models ENS (blue), HRES (red) and AIFS (green) over the forecast period, averaged over all observation sites. Note that the raw models have 6-hour lead time frequency, while the post-processed outputs have 1-hour frequency.
@@
IMPROVER produces probabilistic outputs individually for both deterministic and ensemble input forecasts. For evaluating their accuracy, we use the continuous ranked probability score (CRPS), which measures the error between the predicted probability distribution and the true outcome. The CRPS is defined as the integral of the squared distance between the forecast distribution and the true outcome. Formally, let $X$ be the random variable being forecast, $t:\to\mathrm{P}(X\leq t)$ the predicted cumulative distribution function for some data point, and $x$ the observed value. Then the CRPS for this observation is
$\int_{t=-\infty}^{t=\infty}(\mathrm{P}(X\leq t)-\mathbbm{1}_{(t\geq x)})^{2}dt$
@@
where $\mathbbm{1}_{(t\geq x)}$ is the indicator function which is equal to 1 if $t\geq x$ and 0 otherwise. Smaller CRPS values indicate a better forecast. When evaluating the CRPS, we take the forecast CDF to be the piecewise-linear function obtained by linearly interpolating the probabilities of exceedence at the predicted thresholds. Figure 9 shows the CRPS of the post-processed models and blends. While in general the blend of the two NWP models performs very similarly to the most accurate individual NWP model, the blend of all three models yields a larger improvment. We used the Diebold-Mariano test to compare the two blends, following the same process as described above for the MSE comparison. The results are shown in Figure 10; as is the case with MSE, we see that the all-model blend is significantly better than the NWP-model blend.
@@
Figure 9: Continuous ranked probability score (CRPS) by lead day for post-processed models ENS (blue dashed), HRES (red dashed), and AIFS (green dashed), and blends (NWP models ENS and HRES, yellow; and all models, pink) over the forecast period. The tabular data for these graphs can be found in Tables 4, 5 and 6 in the Appendix.
@@
Forecast reliability measures how well-calibrated a forecast is. Specifically, for a given threshold $t$, a forecast has good reliability if the conditional probability $p_{o}$ that an observation exceeds $t$, given that the forecast probability is $p_{f}$, is approximately equal to $p_{f}$. Figure 11 shows the reliability at the 12-hour lead time (UTC 0) at selected thresholds, roughly corresponding to the 10th, 50th and 90th percentiles of the overall distribution of observations (including all times of day). Note that only a few thresholds are shown here; IMPROVER calculates probabilities at a much larger set of thresholds as described in Section 3. There is no clear trend in these results: all models and blends have similar performance. The deterministic models arguably appear a little better than the ensemble, and the blended models do not seem to offer much advantage. Reliability is still fairly good later in the forecast period: Figure 16 in the Appendix shows the reliability at lead time 228 hours. It has been shown by Ranjan and Gneiting (2010) that blending probability forecasts does not preserve calibration; that is, even if each individual forecast is perfectly calibrated, in general the blend is not. Therefore, calibration could likely be further improved by recalibrating the blend. Functionality to implement the parametric recalibration proposed by Ranjan and Gneiting has recently been implemented in the IMPROVER software library, but is not currently part of the IMPROVER operational system used for this work.
In Figure 12 we show the distributions of probabilistic forecasts at the 12-hour lead time, which allows us to evaluate the forecasts’ sharpness. A sharp forecast is indicated by a deep U-shaped distribution graph, with many values close to 0 or 1, while a flatter graph indicates a less-confident forecast. The 3-model blend is in general less sharp than the 2-model blend, which is in turn less sharp than the sharpest NWP model, namely ENS. Again, this is expected for blended forecasts, and could be corrected with recalibration.
Figure 11: Reliability at the 12-hour lead time for post-processed models ENS (blue dashed), HRES (red dashed), and AIFS (green dashed), and blends (NWP models ENS and HRES, yellow; and all models, pink). Forecast probabilities are binned into 10 equal-width bins, and, for each bin, the average forecast and observed probability of exceeding the threshold is calculated. These values are linearly interpolated to produce the reliability curve. Bins having fewer than 10 data points are not plotted.
Figure 12: Distribution of binned probability forecasts at the 12-hour lead time for post-processed models ENS (blue dashed), HRES (red dashed), and AIFS (green dashed), and blends (NWP models ENS and HRES, yellow; and all models, pink). Forecasts have been binned into 10 equal-width bins. The $y$-axis uses the symmetric log scale, which is linear for values between 0 and 1, and logarithmic for larger values.
@@ 5 Discussion
This study demonstrates that established statistical post-processing methods, originally designed for traditional Numerical Weather Prediction (NWP) models, are also effective when applied to forecasts from the Artificial Intelligence Forecasting System (AIFS). Using the Australian Bureau of Meteorology’s operational IMPROVER system without modification, we observe comparable accuracy improvements in AIFS forecasts to those seen with NWP models. Notably, blending AIFS with traditional NWP models yields consistent gains in forecast skill for both expected value and probabilistic forecasts. These results indicate that AI-based forecasts can be integrated into operational systems using existing tools and workflows, thereby expanding the utility of these forecasts without the need for bespoke post-processing solutions.
A primary finding of this work is that statistical calibration via IMPROVER substantially improves the accuracy of AIFS forecasts. The gains are comparable to those achieved when calibrating traditional models, and in some cases, calibration adds upwards of a day of forecast skill. This reinforces our view that AI models, like their physics-based counterparts, benefit from correction of systematic errors using historical forecast performance. Importantly, the post-processing workflow required no changes, suggesting a high degree of generality in these methods.
The calibrated AIFS output also produces high-quality probabilistic forecasts. For temperature and dew point in particular, the post-processed AIFS shows CRPS values similar to those of the ensemble system (ENS) during the early part of the forecast period. This is notable given that AIFS is a deterministic model, and the result highlights the value of fuzzy thresholding and reliability calibration in extracting probabilistic information from deterministic model inputs.
@@
A limitation of the present work is that the verification period is quite short and does not represent all seasons. It is likely that the input models have seasonally-varying strengths and weaknesses, so it would be valuable to perform the analysis on a longer period. Future work could also consider extending these analyses to AIFS-CRPS (Lang et al., 2024b), an ensemble version of AIFS trained to minimise CRPS. This may further improve the skill of blended forecasts. Additionally, recalibrating the probabilistic blends – particularly in light of known issues with calibration preservation during blending – may yield further improvements.
@@ Acknowledgements.
Figure 13: Mean squared error by lead day for all post-processed NWP models and blends. Note that the BLEND ALL and BLEND ENS AIFS lines are almost identical.
Table 1: Temperature mean squared error by lead day, $(^{\circ}\mathrm{C})^{2}$, corresponding to Figure 6 and Figure 13.
lead day 0 1 2 3 4 5 6 7 8 9
forecast
ECMWF ENS 2.43 2.52 2.73 3.00 3.37 3.86 4.46 5.24 5.99 6.72
ECMWF HRES 2.43 2.61 2.87 3.23 3.73 4.48 5.38 6.41 7.80 9.11
ECMWF AIFS 2.23 2.40 2.58 2.82 3.16 3.75 4.68 5.61 6.51 7.65
BLEND NWP 2.39 2.51 2.71 2.99 3.36 3.85 4.45 5.23 6.00 6.74
BLEND ALL 2.17 2.30 2.47 2.69 3.02 3.53 4.20 4.96 5.71 6.57
BLEND HRES AIFS 2.17 2.31 2.47 2.70 3.03 3.59 4.38 5.24 6.25 7.49
BLEND ENS AIFS 2.19 2.32 2.48 2.70 3.02 3.53 4.21 4.96 5.70 6.56
Table 2: Dew point mean squared error by lead day, $(^{\circ}\mathrm{C})^{2}$, corresponding to Figure 6 and Figure 13.
lead day 0 1 2 3 4 5 6 7 8 9
forecast
ECMWF ENS 2.73 3.11 3.54 4.09 4.88 5.91 7.30 8.89 10.62 12.44
ECMWF HRES 2.82 3.36 3.91 4.69 5.71 7.11 9.24 11.35 13.59 15.86
ECMWF AIFS 2.43 2.88 3.27 3.80 4.54 5.74 7.49 9.48 11.26 13.48
BLEND NWP 2.71 3.10 3.53 4.09 4.86 5.88 7.30 8.91 10.65 12.47
BLEND ALL 2.37 2.74 3.09 3.57 4.26 5.25 6.64 8.28 9.88 11.75
BLEND HRES AIFS 2.37 2.76 3.11 3.61 4.34 5.43 7.09 9.01 10.85 13.03
BLEND ENS AIFS 2.38 2.75 3.09 3.57 4.25 5.25 6.62 8.27 9.88 11.75
Table 3: Wind speed mean squared error by lead day, $(\mathrm{m/s})^{2}$, corresponding to Figure 6 and Figure 13.
lead day 0 1 2 3 4 5 6 7 8 9
forecast
ECMWF ENS 2.77 2.83 2.94 3.17 3.46 3.70 3.96 4.25 4.45 4.64
ECMWF HRES 2.75 2.88 3.02 3.31 3.68 4.10 4.50 4.88 5.22 5.91
ECMWF AIFS 2.74 2.83 2.95 3.14 3.45 3.82 4.23 4.68 5.19 5.54
BLEND NWP 2.74 2.82 2.93 3.15 3.44 3.69 3.95 4.25 4.45 4.64
BLEND ALL 2.68 2.75 2.85 3.05 3.32 3.61 3.89 4.21 4.45 4.61
BLEND HRES AIFS 2.68 2.75 2.87 3.07 3.36 3.73 4.12 4.56 5.03 5.42
BLEND ENS AIFS 2.69 2.75 2.86 3.05 3.32 3.61 3.89 4.18 4.42 4.60
Figure 14: Bias of MSAS analysis against site observations for each hour of the day in UTC time (positive values indicate that MSAS is higher). When comparing with Figure 8, note that the forecast basetime is at UTC 12, so lead time 0 in Figure 8 corresponds to hour 12 in this figure.
Figure 15: Continuous ranked probability score by lead day for all post-processed NWP models and blends. Note that the BLEND ALL and BLEND ENS AIFS lines are almost identical.
Table 4: Temperature continuous ranked probability score by lead day, $^{\circ}\mathrm{C}$, corresponding to Figure 9 and Figure 15.
lead day 0 1 2 3 4 5 6 7 8 9
forecast
ECMWF ENS 0.90 0.94 0.96 1.01 1.06 1.13 1.23 1.33 1.41 1.50
ECMWF HRES 0.87 0.92 0.96 1.02 1.09 1.19 1.31 1.43 1.57 1.70
ECMWF AIFS 0.88 0.92 0.95 0.99 1.04 1.12 1.24 1.35 1.45 1.57
BLEND NWP 0.86 0.91 0.94 0.99 1.05 1.12 1.22 1.32 1.41 1.49
BLEND ALL 0.84 0.88 0.91 0.95 1.00 1.07 1.18 1.27 1.36 1.46
BLEND HRES AIFS 0.84 0.88 0.92 0.96 1.01 1.09 1.20 1.31 1.42 1.55
BLEND ENS AIFS 0.85 0.89 0.92 0.95 1.00 1.07 1.18 1.27 1.36 1.46
Table 5: Dew point continuous ranked probability score by lead day, $^{\circ}\mathrm{C}$, corresponding to Figure 9 and Figure 15.
lead day 0 1 2 3 4 5 6 7 8 9
forecast
ECMWF ENS 0.86 0.92 0.99 1.06 1.16 1.28 1.43 1.58 1.73 1.88
ECMWF HRES 0.87 0.95 1.03 1.13 1.25 1.40 1.60 1.78 1.98 2.15
ECMWF AIFS 0.84 0.91 0.96 1.04 1.13 1.27 1.45 1.64 1.79 1.96
BLEND NWP 0.84 0.91 0.98 1.05 1.15 1.27 1.43 1.58 1.73 1.88
BLEND ALL 0.81 0.87 0.92 0.99 1.08 1.21 1.35 1.51 1.66 1.81
BLEND HRES AIFS 0.82 0.88 0.94 1.01 1.11 1.24 1.41 1.60 1.76 1.93
BLEND ENS AIFS 0.80 0.86 0.92 0.99 1.08 1.20 1.35 1.51 1.66 1.81
Table 6: Wind speed continuous ranked probability score by lead day, $\mathrm{m/s}$, corresponding to Figure 9 and Figure 15.
lead day 0 1 2 3 4 5 6 7 8 9
forecast
ECMWF ENS 0.92 0.93 0.94 0.97 1.00 1.04 1.07 1.11 1.13 1.16
ECMWF HRES 0.89 0.91 0.93 0.97 1.02 1.07 1.12 1.16 1.20 1.28
ECMWF AIFS 0.91 0.92 0.94 0.96 1.00 1.04 1.09 1.14 1.20 1.25
BLEND NWP 0.89 0.91 0.92 0.95 0.99 1.03 1.06 1.10 1.12 1.15
BLEND ALL 0.88 0.90 0.91 0.94 0.97 1.01 1.05 1.09 1.12 1.14
BLEND HRES AIFS 0.88 0.90 0.91 0.94 0.98 1.02 1.07 1.12 1.17 1.22
BLEND ENS AIFS 0.89 0.90 0.92 0.94 0.97 1.01 1.05 1.09 1.12 1.14
Figure 16: Reliability at the 228-hour lead time for post-processed models ENS (blue dashed), HRES (red dashed), and AIFS (green dashed), and blends (NWP models ENS and HRES, yellow; and all models, pink). Forecast probabilities are binned into 10 equal-width bins, and, for each bin, the average forecast and observed probability of exceeding the threshold is calculated. These values are linearly interpolated to produce the reliability curve. Bins having fewer than 10 data points are not plotted.